# Adapting Web Agents with Synthetic Supervision

### v1

@@ Abstract
Web agents struggle to adapt to new websites due to the scarcity of environment specific tasks and demonstrations. Recent works have explored synthetic data generation to address this challenge, however, they suffer from data quality issues where synthesized tasks contain hallucinations that cannot be executed, and collected trajectories are noisy with redundant or misaligned actions. In this paper, we propose SynthAgent, a fully synthetic supervision framework that aims at improving synthetic data quality via dual refinement of both tasks and trajectories. Our approach begins by synthesizing diverse tasks through categorized exploration of web elements, ensuring efficient coverage of the target environment. During trajectory collection, we refine tasks when conflicts with actual observations are detected, mitigating hallucinations while maintaining task consistency. After collection, we conduct trajectory refinement with a global context to mitigate potential noise or misalignments. Finally, we fine-tune open-source web agents on the refined synthetic data to adapt them to the target environment. Experimental results demonstrate that SynthAgent outperforms existing synthetic data methods, validating the importance of high-quality synthetic supervision. The code will be publicly available at https://github.com/aiming-lab/SynthAgent.
@@ 1 Introduction
Large language models (LLMs) with multimodal capabilities have enabled a new wave of web agents capable of autonomously completing complex tasks on the internet (nakano2021webgpt; hong2024metagpt; he2024openwebvoyager; agashe2025agent; li2025websailor). These agents take user instructions and then interact with websites to accomplish tasks, showing promising results on standardized benchmarks (zhou2024webarenarealisticwebenvironment; yao2022webshop; wei2025browsecomp; deng2023mind2webgeneralistagentweb). However, a persistent challenge is that web agents struggle to adapt to new websites not seen during training (zhou2024webarenarealisticwebenvironment; pahuja2025explorerscalingexplorationdrivenweb; he2024openwebvoyager), because new environments often lack sufficient task demonstrations. And issuing environment specific tasks and collecting trajectories demonstrated by human on every new website can be prohibitively expensive. Existing training datasets are either limited to a few domains or lack diversity (xu2024agenttrek; deng2023mind2webgeneralistagentweb; xie2024osworld; chen2024gui), thus when an agent is deployed on a new website, it could frequently encounter unfamiliar states or tasks for which it has no experience. This raises the challenge of how to effectively adapt web agents to new environments with minimal human involvement.
@@
A straightforward way to improve an agent’s performance on a new website environment is to collect more environment specific training data. However, traditional agentic data collection relies on human experts or manually scripted tasks (zhou2024webarenarealisticwebenvironment; lu2024weblinx; deng2023mind2webgeneralistagentweb; mitra2024agentinstruct). Such approaches are labor-intensive and time consuming which can not easily scale. This may lead to significant gaps between training data and the real-world environments where agents are deployed. Without abundant experience in the new and unfamiliar environment, it is challenging for agents to adapt and perform effectively.
To address the data scarcity issue in web agents, as shown in Figure 1, synthetic data generation (xu2024agenttrek; su2025learn; sun2025osgenesisautomatingguiagent; pahuja2025explorerscalingexplorationdrivenweb) has emerged as a promising solution, which uses LLMs to automatically generate and collect data for training agents. Traditional self-instruct methods (wang2023selfinstructaligninglanguagemodels; alpaca) can be used for generating agentic tasks, while it is hard to directly generate information intensive trajectory which often contains a series of observations such as HTML code and screenshots. Recently, sun2025osgenesisautomatingguiagent propose generating diverse tasks based on the observations of the environment changes, while hallucinations in generated tasks can make it impossible to collect corresponding trajectory. Explorer (pahuja2025explorerscalingexplorationdrivenweb) addresses this issue by iteratively refining tasks based on the agent’s observations during interaction with the environment (i.e., trajectory collection). However, since the initial tasks are often underspecified, the beginning trajectories tend to contain substantial noise. It is challenging to fully synthesize high-quality agentic training data for a new environment without any human involvement.
In this paper, we propose SynthAgent, a synthetic supervision framework for adapting web agents to new environments. Compared to existing approaches, SynthAgent aims at ensuring the quality in synthesis by introducing dual refinement to both task and trajectory. Specifically, our method consists of four stages: (1) First, we synthesize domain specific agentic tasks by extensively exploring the environment and strategically categorizing web elements. (2) Then, we deploy an agent to collect trajectories for the synthesized tasks, while refining tasks based on new observations to mitigate hallucinations. (3) After collection, we further refine the collected trajectories using global context to mitigate potential noise or misalignments. (4) Finally, we fine-tune an open-source web agent on the refined synthetic data to adapt it to the target environment. In contrast to existing works focusing on data synthesis, our approach aims at improving the quality of synthetic data via dual refinement.
@@
    We propose SynthAgent, a fully synthetic supervision framework for adapting web agents to new environments without test task leakage or human involvement.
@@
    We identify critical quality issues in existing synthetic data generation methods, i.e., hallucinations in tasks and noise in trajectories. We introduce dual refinement to substantially improve synthetic data quality.
@@
    Experimental results on controllable web environments demonstrate that SynthAgent outperforms existing synthetic data baselines, validating the importance of high-quality synthetic supervision for web agent adaptation.
@@ 2.1 Web Agent
Recent advances in LLMs have driven interest in developing agents that combine reasoning and interaction (wei2022chain; yao2023tree; wang2022self; lightman2023let; guo2025deepseek). ReAct (yao2023reactsynergizingreasoningacting) introduces the interleaving of reasoning and actions after observations, while following works (gao2023palprogramaidedlanguagemodels; hong2024metagpt; wu2024autogen; nakano2021webgpt) explore the use of tools integration and planning to enhance agent capabilities. However, these works primarily rely on human-annotated data for training, which is costly to scale and difficult to adapt to new environments. Another series of works (lu2025ui; qiu2025alita; wei2025webagent) use reinforcement learning to train web agents, but they often require supervision from expert annotation and are impractical for complex and realistic websites due to the high cost of online training. Using synthetic data to train web agents has also gained traction as a scalable alternative, especially for new environments (wang2024survey; liu2025advances).
@@ 2.2 Data Synthesis
Data synthesis has emerged as a powerful paradigm to address data scarcity across various fields, allowing models to learn from automatically generated examples rather than costly human annotations. Early works such as Self-Instruct (wang2023selfinstructaligninglanguagemodels) and Alpaca (alpaca) leverage advanced but closed-source LLMs to bootstrap instruction tuning data for training smaller open-source models. Meanwhile, other studies (zelikman2022star; wang2023democratizing; ge2024scaling; zhao2025absolute) explore synthesizing large-scale training data to enhance the reasoning performance of smaller LLMs. These successes demonstrate the promising potential of synthetic data for adapting models to new tasks and domains. However, in agent scenarios, many works focus solely on synthesizing trajectories while directly using tasks from the test set to train the model (chen2024agent; zhou2024webarenarealisticwebenvironment; zhang2025symbiotic), raising serious concerns about test set leakage. Fortunately, recent works (he2024openwebvoyager; su2025learn; sun2025osgenesisautomatingguiagent; pahuja2025explorerscalingexplorationdrivenweb) have explored data synthesis for web agents from scratch, which includes both task and trajectory synthesis.
@@ 2.3 Synthetic data for Web Agents
Inspired by the success of distilling knowledge from large LLMs into smaller models (wang2023selfinstructaligninglanguagemodels; alpaca; xu2023wizardlm; wang2023democratizing), synthetic data has been increasingly used to train web agents. he2024openwebvoyager introduce self-instruct for agentic task generation, where it mainly operates on the surface of websites, thus the synthesized tasks are often simple and repetitive. AgentTrek (xu2024agenttrek) collects training data by scraping web tutorials from offline corpus, which are often outdated and not feasible for new environments. WebSynthesis (gao2025websynthesisworldmodelguidedmctsefficient) employs a learned world model of web interfaces to simulate interactions, but its reliance on LLM-based environment can introduce additional hallucinations. Explorer (pahuja2025explorerscalingexplorationdrivenweb) synthesizes tasks and trajectories simultaneously and refines initial tasks based on observations. However, since these tasks are often underspecified at the beginning, the trajectories may contain substantial noise due to wandering. OS-Genesis (sun2025osgenesisautomatingguiagent) collects low-level actions through GUI exploration and infers high-level tasks from them. However, its task proposals rely solely on single-step environment changes, often leading to severe hallucinations. Additionally, OS-Genesis explores environments via random actions, which produces repetitive tasks and inefficient exploration. To address these limitations in data quality and diversity, SynthAgent introduces dual refinement for both task and trajectory synthesis, effectively reducing hallucinations and noise. Further, SynthAgent uses elements categorization to guide exploration and task generation, enhancing both diversity and coverage of the synthesized data.
Figure 2: Overview of SynthAgent and its comparison with baseline methods: OS-Genesis (sun2025osgenesisautomatingguiagent) and Explorer (pahuja2025explorerscalingexplorationdrivenweb). SynthAgent consists of four steps: (1) Task Synthesis with Categorized Exploration, which synthesizes diverse, environment specific tasks through categorized exploration. (2) Task Refinement during Trajectory Collection, which collects trajectories while refining task descriptions based on new observations to mitigate potential hallucinations. (3) Trajectory Refinement, which further improves collected trajectories using global context to edit noisy actions. (4) Agent Fine-tuning, which adapts the web agent to new environments under fully synthetic supervision. The right panel highlights the key advantage of SynthAgent over baselines, i.e., the proposed dual refinement of both task and trajectory synthesis can lead to better synthetic data.
@@ 3.1 Task Synthesis
Following OS-Genesis (sun2025osgenesisautomatingguiagent), we explore the environment to synthesize diverse tasks. It collects interaction triplets $(o_{t},a_{t},o_{t+1})$ by executing an action $a_{t}$ on a randomly selected element of the current page observation $o_{t}$, resulting in a new webpage $o_{t+1}$. Then, it prompts an LLM to propose high-level tasks based on each triplet. However, such random exploration can lead to redundant interactions and limited task diversity.
To address this issue, we propose categorized exploration to systematically cover different types of web elements and interactions. Specifically, at each visited page $o_{t}$ with element set containing text, images, inputs, buttons, links and etc., we classify elements into several categories (e.g., Account Management, Search & Filters, Shopping Content and etc. for a shopping webpage) based on their functional intention. This partition is enabled by prompting an LLM with the page structure and simple heuristics such as element role and name. From each category we uniformly sample up to $2$ unvisited elements to interact with, which can then yield a batch of interaction triplets. We also maintain a URL pool to record newly discovered pages, for resuming exploration to deeper pages later.
For each collected triplet $(o_{t},a_{t},o_{t+1})$, we prompt the LLM to propose a high-level task $\tau$ describing user’s intention that (1) is achievable from $o_{t}$ by plausible multi-step interactions and (2) is highly possible to be completed in $o_{t+1}$ or subsequent pages. The LLM is requested to imagine a broader objective that requires multiple steps to accomplish, where the single action $a_{t}$ is just one step towards the goal. These synthesized tasks are thus more environment-aware and diverse, covering various categories and depths of target website.
@@ 3.2 Task Refinement
After task proposal, OS-Genesis deploys the agent to interact with the environment to collect trajectories for each synthesized task. Note that its proposed task is based solely on a single interaction, thus it often contains hallucinations (e.g., assuming an option or account state that does not exist), causing agent cannot complete the task. Explorer (pahuja2025explorerscalingexplorationdrivenweb) combines task proposal and trajectory collection into one stage. It first proposes a coarse task from the homepage and then continuously refines the task based on new observations during trajectory collection. This design implicitly acknowledges that the initial task is underspecified which means following refinement is more about filling in details rather than correcting hallucinations. Empirically, we observe that Explorer frequently changes the task intent during execution ($8.6$ times vs. $2.0$ times of SynthAgent), which often leads to overly long trajectories that fail to complete within the budget ($68.3\%$ samples vs. $6.3\%$ samples of SynthAgent). To mitigate hallucinations in task synthesis, we separate task proposal and trajectory collection into two distinct steps, in order to avoid underspecified initial tasks. Building on our categorized exploration, the initial tasks are already more detailed, while it still may contain hallucinations due to limited observations. Therefore, we introduce the following task refinement during trajectory collection to reduce such noise.
@@
Let $\tau_{t}$ denote the task specification at time $t$ and $h_{t}=(o_{1},a_{1},\ldots,o_{t})$ the partial trace. The task is likely to contain hallucinations if it conflicts with the observation on the realistic environment. Thus, we trigger refinement when detect a conflict with a lightweight predicate $\mathcal{C}(h_{t},\tau_{t}){=}\text{true}$, where
@@
Upon a trigger, we call the LLM to follow four evidence-driven principles to refine the task: (1) concretize missing details, (2) align with actual observations, (3) downscope or simplify when blocked, and (4) preserve similar task category. The replacement task $\tau_{t+1}$ is expected to resolve the conflict while remaining as close as possible to the original intent. The following actions are then taken to continue trajectory collection, in order to complete the new task $\tau_{t+1}$. After execution, the last specification is the final task $\tau^{\star}$ along with the collected trajectory $h_{T}$. Note that refinement processes can introduce discontinuity and noise in the trajectory, which we address in the next step.

### v2

@@ Abstract
Web agents struggle to adapt to new websites due to the scarcity of environment specific tasks and demonstrations. Recent works have explored synthetic data generation to address this challenge, however, they suffer from data quality issues where synthesized tasks contain hallucinations that cannot be executed, and collected trajectories are noisy with redundant or misaligned actions. In this paper, we propose SynthAgent, a fully synthetic supervision framework that aims at improving synthetic data quality via dual refinement of both tasks and trajectories. Our approach begins by synthesizing diverse tasks through categorized exploration of web elements, ensuring efficient coverage of the target environment. During trajectory collection, tasks are refined only when conflicts with observations are detected, which mitigates hallucinations while preserving task consistency. After collection, we conduct trajectory refinement with global context to mitigate potential noise or misalignments. Finally, we fine-tune open-source web agents on the refined synthetic data to adapt them to the target environment. Experimental results demonstrate that SynthAgent outperforms existing synthetic data methods, validating the importance of high-quality synthetic supervision. The code is publicly available at https://github.com/aiming-lab/SynthAgent.
@@ 1 Introduction
Large language models (LLMs) with multimodal capabilities have enabled a new wave of web agents capable of autonomously completing complex tasks on the internet (hong2024metagpt; he2024openwebvoyager; agashe2025agent; li2025websailor; yang2025agentoccam). These agents take user instructions and then interact with websites to accomplish tasks, showing promising results on standardized benchmarks (zhou2024webarenarealisticwebenvironment; yao2022webshop; wei2025browsecomp; deng2023mind2webgeneralistagentweb). However, a persistent challenge is that web agents struggle to adapt to new websites not seen during training (zhou2024webarenarealisticwebenvironment; pahuja2025explorerscalingexplorationdrivenweb; he2024openwebvoyager), because new environments often lack sufficient task demonstrations. And issuing environment specific tasks and collecting trajectories by human on every new website can be expensive. Existing training datasets are either limited to a few domains or lack diversity (xu2024agenttrek; deng2023mind2webgeneralistagentweb; xie2024osworld; chen2024gui), thus when an agent is deployed on a new website, it could frequently encounter unfamiliar states or tasks for which it has no experience. This raises the challenge of how to effectively adapt web agents to new environments without human involvement.
@@
A straightforward way to improve an agent’s performance on a new website environment is to collect more environment specific training data. However, traditional agentic data collection relies on human experts or manually scripted tasks (zhou2024webarenarealisticwebenvironment; lu2024weblinx; deng2023mind2webgeneralistagentweb; mitra2024agentinstruct). Such approaches are labor-intensive and time consuming, which cannot be easily scaled. This may lead to significant gaps between training data and the real-world environments where agents are deployed. Without abundant experience in the new and unfamiliar environment, it is challenging for agents to adapt and perform effectively.
To address the data scarcity issue, as shown in Figure 1, synthetic data (xu2024agenttrek; su2025learn; sun2025osgenesisautomatingguiagent; pahuja2025explorerscalingexplorationdrivenweb) has emerged as a promising solution, which uses LLMs to generate and collect data for training. However, we identify two critical limitations in existing web agent data generation pipelines that prevent scaling high-quality synthetic data: (1) Task hallucinations: methods like OS-Genesis (sun2025osgenesisautomatingguiagent) propose tasks from single-step observations, often generating goals that reference non-existent elements or impossible states. (2) Trajectory noise: methods like Explorer (pahuja2025explorerscalingexplorationdrivenweb) propose a coarse task from the homepage and then continuously refine underspecified tasks during execution, causing agent wandering and producing trajectories where early segments are misaligned with the final task. These issues are not merely quality problems but stem from a design tension: task quality requires environment grounding, yet grounding during execution introduces trajectory noise.
In this paper, we propose SynthAgent, a fully synthetic supervision framework that resolves this tension through dual refinement: tasks are refined during execution to mitigate task hallucinations, and trajectories are subsequently refined to mitigate potential noise introduced by task edits. Specifically, our method consists of four stages: (1) First, we synthesize diverse tasks through the proposed categorized exploration that systematically covers functional groups of web elements, improving both task diversity and exploration coverage. (2) Then, during trajectory collection, we apply task refinement triggered by explicit conflict detection, correcting hallucinations while minimizing unnecessary task changes. (3) After collection, we conduct trajectory refinement with global context, mitigating noise and misalignments introduced by task edits or agent wandering. (4) Finally, we fine-tune open-source web agents on the refined synthetic data. The key insight is that task and trajectory refinement are synergistic: task refinement enables feasible execution but introduces noise, which trajectory refinement subsequently mitigates.
@@
    We propose SynthAgent, a fully synthetic supervision framework for effectively adapting web agents to new environments without test task leakage or human involvement.
@@
    We identify a critical tension in existing methods: environment grounding during task execution inevitably introduces trajectory noise. We resolve this through a dual refinement strategy. We also propose categorized exploration to enhance task diversity and coverage.
@@
    Extensive experiments show that SynthAgent substantially outperforms existing synthetic data methods across diverse websites, highlighting the effectiveness of high-quality synthetic supervision in adapting web agents.
@@ 2.1 Web Agent
Recent advances in LLMs have driven interest in developing agents that combine reasoning and interaction (wei2022chain; yao2023tree; wang2022self; lightman2023let; guo2025deepseek). ReAct (yao2023reactsynergizingreasoningacting) introduces the interleaving of reasoning and actions after observations, while following works (gao2023palprogramaidedlanguagemodels; hong2024metagpt; wu2024autogen; nakano2021webgpt; yang2025agentoccam) explore tools integration, planning, and observation-action alignment to enhance agent capabilities. However, these typically rely on human annotation, which is costly to scale and adapt to new environments. Another series of works (lu2025ui; qiu2025alita; wei2025webagent) use reinforcement learning to train web agents, but they often require supervision from annotation and are impractical for complex and realistic websites due to the cost of online training. Using synthetic data to train web agents has gained traction as a scalable alternative, especially for new environments (wang2024survey; liu2025advances).
@@ 2.2 Data Synthesis
Data synthesis has emerged as a powerful paradigm for addressing data scarcity across various fields, allowing models to learn from automatically generated examples rather than costly human annotations. Early works such as Self-Instruct (wang2023selfinstructaligninglanguagemodels) and Alpaca (alpaca) leverage advanced but closed-source LLMs to bootstrap instruction tuning data to train smaller open-source models. Meanwhile, other studies (zelikman2022star; wang2023democratizing; ge2024scaling; zhao2025absolute) explore synthesizing large-scale training data to enhance the reasoning performance of smaller LLMs. These successes demonstrate the promising potential of synthetic data for adapting models to new tasks and domains. However, in agent scenarios, many works focus solely on synthesizing trajectories while directly using tasks from the test set to train the model (chen2024agent; zhou2024webarenarealisticwebenvironment; zhang2025symbiotic), raising serious concerns about test set leakage. Fortunately, recent works (he2024openwebvoyager; su2025learn; sun2025osgenesisautomatingguiagent; pahuja2025explorerscalingexplorationdrivenweb) have explored data synthesis for web agents from scratch, including both task and trajectory synthesis.
@@ 2.3 Synthetic data for Web Agents
Synthetic data has been increasingly used to train web agents (wang2023selfinstructaligninglanguagemodels; alpaca; xu2023wizardlm; wang2023democratizing). he2024openwebvoyager introduce self-instruct for agentic task generation, where it mainly operates on the surface of websites, thus the synthesized tasks are often simple and repetitive. AgentTrek (xu2024agenttrek) collects training data by scraping web tutorials from offline corpus, which can be outdated and not feasible for new environments. Synatra (ou2024synatra) generates offline HTML trajectories by injecting web knowledge from external sources (e.g., wikiHow) without interacting with real websites, which limits the grounding of synthesized data. NNetNav (murty2025nnetnavunsupervisedlearningbrowser) employs an LLM-based trajectory labeler to prune exploration based on instruction hierarchies, but does not address data quality issue. WebSynthesis (gao2025websynthesisworldmodelguidedmctsefficient) employs a learned world model of web interfaces to simulate interactions, but its LLM-based environment can introduce additional hallucinations.
Two recent works are most related to ours. Explorer (pahuja2025explorerscalingexplorationdrivenweb) refines underspecified initial tasks during execution, but produces noisy trajectories due to wandering and endless refinement. OS-Genesis (sun2025osgenesisautomatingguiagent) pioneered reverse task synthesis through random GUI exploration, but suffers from hallucinations in single-step-based task proposals and inefficient coverage. In contrast, SynthAgent introduces structural improvements: (1) categorized exploration for systematic coverage and improved task diversity, (2) task refinement that detects and corrects hallucinations based on actual observations, and (3) trajectory refinement with global context that removes noise from task edits or agent wandering.
Figure 2: Overview of SynthAgent compared to baseline methods OS-Genesis (sun2025osgenesisautomatingguiagent) and Explorer (pahuja2025explorerscalingexplorationdrivenweb). SynthAgent consists of four steps: (1) Task Synthesis, generating diverse tasks via categorized exploration; (2) Task Refinement, refining tasks during trajectory collection to avoid task hallucinations; (3) Trajectory Refinement, globally refining collected trajectories to remove noisy actions; and (4) Agent Fine-tuning, adapting web agents using fully synthetic data. The right panel emphasizes the advantage of our dual refinement strategy over baseline methods, leading to higher-quality synthetic web agent data.
@@ 3.1 Task Synthesis
A key challenge in task synthesis is achieving both diversity (covering different functionalities) and feasibility (avoiding hallucinated goals). Prior work explores environments via random interactions with the environment (sun2025osgenesisautomatingguiagent), which produces redundant interactions and limited coverage. We address this through categorized exploration that systematically partitions web elements by their functional roles. Specifically, at each visited page $o_{t}$, we classify interactive elements (buttons, inputs, links, etc.) with their names and roles into functional categories (e.g., Account Management, Search & Filters, Shopping Content) by prompting an LLM with the page structure. From each category, we uniformly sample up to $2$ unvisited elements to interact with, collecting interaction triplets $(o_{t},a_{t},o_{t+1})$ where $a_{t}$ is the corresponding action for the sampled element and $o_{t+1}$ is the resulting page. We maintain a URL pool to track discovered pages for deeper exploration.
For each triplet, we prompt the LLM to propose a high-level task $\tau$ that (1) is achievable from $o_{t}$ through multi-step interactions and (2) is likely completable in $o_{t+1}$ or subsequent pages. The LLM imagines a broader objective where $a_{t}$ is one step towards the goal. This design ensures synthesized tasks are environment-grounded and diverse, covering various functional categories and interaction depths. Statistics show our exploration yields an average of $6$ functional categories per page with stable granularity across different websites. See detailed statistics in Appendix A.2.
@@ 3.2 Task Refinement
After task proposal, OS-Genesis deploys the agent to interact with the environment to collect trajectories for each synthesized task. Note that its proposed task is based solely on a single interaction, thus it often contains hallucinations (e.g., assuming an option or account state that does not exist), causing the agent to fail to complete the task.
For our method, even with categorized exploration, synthesized tasks may contain hallucinations due to limited observations. A natural solution is to refine tasks during execution when conflicts arise. However, how refinement is triggered matters critically: Explorer (pahuja2025explorerscalingexplorationdrivenweb) combines task proposal and trajectory collection into one stage. It first proposes a coarse task from the homepage and then refines the task continuously at every step. This design implicitly acknowledges that the initial task is underspecified, meaning the following refinement is more about filling in details rather than correcting hallucinations. Empirically, we observe that Explorer frequently changes the task intent during execution ($8.6$ times vs. $2.0$ times of SynthAgent), which often leads to overly long trajectories that fail to complete within the budget ($68.3\%$ samples vs. $6.3\%$ samples of SynthAgent). In contrast, we design conflict-triggered refinement that activates only when explicit mismatches are detected, resulting in only $2.0$ refinements on average and $6.3\%$ budget exceedance. The key difference is that our initial tasks are already well-specified, so the refinement should focus on correcting hallucinations rather than continuously changing intention.
@@
Let $\tau_{t}$ denote the task specification at time $t$ and $h_{t}=(o_{1},a_{1},\ldots,o_{t})$ the partial trace. The task is likely to contain hallucinations if it conflicts with the observation on the realistic environment. Thus, we trigger refinement when we detect a conflict with a lightweight predicate $\mathcal{C}(h_{t},\tau_{t}){=}\text{true}$, where
@@
Upon a trigger, we call the LLM to follow four evidence-driven principles to refine the task: (1) concretize missing details, (2) align with actual observations, (3) downscope or simplify when blocked, and (4) preserve a similar task category. The replacement task $\tau_{t+1}$ is expected to resolve the conflict while remaining as close as possible to the original intent. The following actions are then taken to continue trajectory collection in order to complete the new task $\tau_{t+1}$. After execution, the final specification is the task $\tau^{\star}$ along with the collected trajectory $h_{T}$. Note that the refinement process can introduce discontinuities and noise into the trajectory, which we address in the next step.